### v1

# SHuBERT: Self-Supervised Sign Language Representation Learning via Multi-Stream Cluster Prediction

Abstract
Sign language processing has traditionally relied on task-specific models, limiting the potential for transfer learning across tasks. We introduce SHuBERT (Sign Hidden-Unit BERT), a self-supervised transformer encoder that learns strong representations from approximately 1,000 hours of American Sign Language (ASL) video content. Inspired by the success of the HuBERT speech representation model, SHuBERT adapts masked prediction for multi-stream visual sign language input, learning to predict multiple targets for corresponding to clustered hand, face, and body pose streams. SHuBERT achieves state-of-the-art performance across multiple benchmarks. On sign language translation, it outperforms prior methods trained on publicly available data on the How2Sign (+0.7 BLEU), OpenASL (+10.0 BLEU), and FLEURS-ASL (+0.3 BLEU) benchmarks. Similarly for isolated sign language recognition, SHuBERT’s accuracy surpasses that of specialized models on ASL-Citizen (+5%) and SEM-LEX (+20.6%), while coming close to them on WLASL2000 (-3%). Ablation studies confirm the contribution of each component of the approach.
1 Introduction
Sign languages are used by millions of individuals worldwide, but the development of automatic sign language processing systems lags far behind that of systems for spoken and written language [62]. Unlike speech or text processing, sign language presents unique challenges due to its multi-channel nature, combining manual gestures, facial expressions, and body movements, which can be quick and highly coarticulated [3].
Existing approaches to sign language processing have typically relied on models designed and trained for specific tasks, such as translation from signed to written languages [6, 33, 63], isolated sign recognition [27, 25], and fingerspelling detection and recognition [49, 14, 17]. These approaches often require large amounts of annotated data, which is particularly scarce for sign languages. This limitation has historically constrained the performance and scalability of sign language processing systems.
Figure 1: Comparison between our results using fine-tuned SHuBERT and results of the previous state-of-the-art task-specific models on a suite of multiple tasks, datasets, and metrics. Note: The orange shade does not represent a single model but a collection of the previous SOTA results on these tasks, datasets, and metrics. The SHuBERT-based results improve on all but one task-specific SOTA model in our experiments. See Sec. 4 for more experimental details.
Figure 2: SHuBERT pre-training. (a) We locate a set of landmarks in each frame of the input video using MediaPipe [35], with inter-frame interpolation to fill in missing landmarks. From these, we extract the upper body pose, crop the hand and face regions, and blur and partially mask the face crop is blurred for a measure of privacy. (b) We extract body pose and DINOv2 features for the hands and face, yielding a four-stream representation (two hands, face, body pose) for each frame. (c) We assign the feature vectors for frame $t$ to cluster indices using pre-computed $k$-means clusters, yielding assignments $(f_{t},l_{t},r_{t},b_{t})\in[k]^{4}$ for face, left and right hand and body pose, respectively. (d) We partially mask the features, which form the input to a transformer encoder. The output of the transformer is fed to a linear classifier (e) predicting the cluster assignments for each frame, $(\widehat{f}_{t},\widehat{l}_{t},\widehat{r}_{t},\widehat{b}_{t})$. We train the linear layer and the encoder using the log-loss of the predicted vs. true assignments for the masked tokens.
Recent advances in self-supervised learning of representations for natural language and speech have produced representations that perform well across a wide range of tasks. For example, encoder models like BERT for written language [12, 31] and HuBERT for speech [20] are learned from unlabeled data by predicting masked portions of the input. Self-supervised learning seems particularly relevant for sign languages, for which annotated data are particularly scarce. But the unique visual properties of sign languages, combining interacting fine-grained hand and facial gestures and global body motion, demand a specialized approach.
In this work, we present SHuBERT (Sign Hidden-Unit BERT) (see  Fig. 2), a self-supervised representation learning approach that adapts the masked prediction paradigm to the characteristics of sign language video. SHuBERT learns contextual, transformer-based representations of video frames by predicting cluster assignments of multiple masked video streams of the hands, face, and body pose features. We demonstrate that the resulting learned representations transfer effectively to multiple downstream tasks, achieving state-of-the-art performance on several sign language translation (SLT) benchmarks, as well as on multiple isolated sign recognition (ISLR) benchmarks. Our results suggest that a self-supervised learning approach to sign language representation learning is not only possible, but can outperform specialized architectures while requiring less task-specific annotation. We conclude that SHuBERT is a promising candidate for the role of foundation model [4] for sign language video processing.
2 Related Work
Advances in sign language processing have often been based on inspiration from methods found to be successful for other video domains and for text. This section first examines influential pre-training methods in text and speech processing that inform SHuBERT, and then explores specific developments in sign language understanding, including both supervised and self-supervised approaches.
2.1 Pre-Training for Text And Speech
Pre-training has become a cornerstone of modern language processing across modalities. In natural language (text) processing, encoder models like BERT [12] and its variants (e.g., RoBERTa [34] and ALBERT [31]), based on masked language modeling, have served as dependable representations that can be used and adapted for a variety of language understanding tasks [4]. Pre-trained encoder-decoder models, like T5 [43], unify multiple text-to-text tasks in a single model. Finally, decoder-only models like the GPT series [42] have demonstrated the power of auto-regressive pre-training, when very large-scale training data is available and the downstream tasks can be framed in terms of auto-regressive generation.
In speech processing, self-supervised learning approaches have taken inspiration from text encoder models while addressing the unique challenges posed by continuous audio signals, which have no inherent segmentation into tokens nor a pre-defined token vocabulary. For example, Hidden-Unit BERT (HuBERT) [20] adapted the objective of BERT by adding an offline clustering step to provide pseudo-labels for masked prediction. Other approaches include wav2vec 2.0 [2], which uses a contrastive objective, and AV-HuBERT [51] which extends HuBERT to audiovisual speech (i.e., speech along with video of the speaker’s mouth). Signed languages share similar challenges to speech, with no pre-existing token lexicon and variable-length units (gestures) with no explicit boundaries, and our approach takes inspiration from HuBERT.
2.2 Pre-Training for Sign Language Understanding
Sign language processing research has explored both supervised and self-supervised pre-training approaches for individual tasks.
For translation of sign language video, the supervised pre-training approach has focused on first (pre-)training a translation model on a large (but often noisy) out-of-domain dataset, followed by fine-tuning on a smaller in-domain dataset. This type of pre-training leverages large collections of annotated data, with some systems [55, 54, 52, 63] trained on up to 6,600 hours of sign language content [55] to achieve state-of-the-art performance. However, much of this data remains private, including the largest dataset used by [63]. In the self-supervised line of work, SSVP-SLT [44] adapts masked autoencoders for sign language video, achieving competitive performance on the sign language translation task. These approaches, while effective, often involve substantial computational resources: the supervised model of [63] was trained on 128 TPU-v3 chips for 20 days, while the MAE-like self-supervised approach of [44] used 64 A100 GPUs for 14 days.
Other lines of work (e.g., [8]) have use a pre-trained S3D model [60], which requires the video sequence to be segmented into chunks, with each chunk treated as independent by the model. This form of pre-training therefore doesn’t capture long temporal context. SHuBERT instead learns contextual frame representations and can therefore operate on long video sequences directly.
For isolated sign language recognition, several self-supervised approaches have shown promise. SignBERT [22] is a self-supervised model that treats hand poses as visual tokens and learns via masked prediction of hand joints. An extension of this approach is SignBERT+ [23], which is a self-supervised pre-training approach for sign language that focuses on hand pose modeling with a specialized hand-model-aware decoder. In contrast, SHuBERT models all components of sign language by treating it as a multi-stream signal, simultaneously modeling the face, hands, and body pose. Another recent study compared multiple self-supervised techniques for ISLR, finding masked autoencoders (MAE) particularly effective [45].
2.3 Multi-Stream Models of Sign Language
Another recent translation approach, SignMusketeers [19], found that it is possible to achieve similar performance to SSVP-SLT while using less than 5% of the compute, by using a multi-stream model of the hands, face, and body. SHuBERT also adopts the idea of multiple streams. However, unlike prior multi-stream approaches, SHuBERT learns a self-supervised representation that learns from the multiple streams jointly and that performs well on multiple tasks. Several previous methods have taken advantage of the observation that sign language naturally decomposes into multiple streams of hand, face, and body motions. In addition to the aforementioned SignMusketeers, prior work has developed multi-stream models for fingerspelling recognition (which combines hand and mouthing gestures) [48], sign language translation  [50, 7, 65, 9], and isolated sign language recognition [41, 24]. SHuBERT also uses a decomposition into multiple streams of information, but learns a representation in a self-supervised way that combines information from all of them.
Figure 3: Sample frames of several signers and the corresponding input channels used for SHuBERT: blurred face crop, left hand crop, and right hand crop. In addition to the face and left/right hand features extracted from these crops, each frame is represented by an additional feature vector corresponding to the upper body pose extracted from MediaPipe (see Fig. 2).
3 Sign Hidden-Unit BERT (SHuBERT)
SHuBERT is a transformer encoder [56] that learns contextualized representations of sign language video frames through self-supervised learning. The pre-training approach is outlined in Fig. 2. In the following sections we describe the video features used in SHuBERT (Sec. 3.1) and the self-supervised training approach (Sec. 3.2).
3.1 Multi-Stream Feature Pre-Processing
Fig. 3 provides examples of SHuBERT’s input video features, described in detail below.
Handshapes
We use the MediaPipe Hand Landmarker [^1] https://ai.google.dev/edge/mediapipe/solutions/vision/hand_landmarker model, which has hand detection accuracy 95%. Upon inspection, we find that the majority of the remaining 5% of “failed” detections occur when the hands are outside the frame. For these cases, we interpolate from the nearest frames with successful detections. Dilated bounding boxes for the detected hand landmarks (for both left and right hand) are cropped and resized to 224×224.
Facial Features
The signer’s face contains important non-manual markers for sign languages [5]. Previous approaches either use the full face, compromising privacy [19], or completely blur the whole face in an attempt to protect privacy, potentially losing essential non-manual markers [44]. Our design attempts to balance the need to preserve linguistic information conveyed by the face with the goal of enhancing the signer’s privacy. We identify the whole face, mouth and eye regions in the frame from the relevant MediaPipe facial landmarks. The face pixels are greyed out except for the pixels in the mouth and eye regions. We then apply Gaussian blur to the entire face region, resize it to 224×224, and feed it into our DINO-based extractor.
Self-Supervised Image Feature Extractor
For face and hand image crops, while most works use pose, we use DINOv2 [38] for feature extraction, which has proven to be a robust feature extractor for sign language in previous work [59, 19]. An additional benefit of DINOv2 representations is that they yield meaningful clusters when quantized [64], an important property since SHuBERT training targets are clustered input features. While most other works use keypoints estimation tools, it has shown to have some weakness to adequately capture sign language handshapes [36] and facial expressions [30].
To adapt the general image feature extractor to a face feature extractor for sign language, we use 5 million random face crops from videos in the YouTube-ASL [55] dataset to fine-tune DINOv2 for 1 epoch. We do the same for the hand feature extractor, using 5 million random hands (mixing left and right hand crops) from YouTube-ASL [55]. For the face and hand streams, crop regions of interest (ROIs) are processed through the face or hand fine-tuned DINOv2 models, yielding a 384-dimensional feature vector per crop.
Body Posture
For body pose, we extract seven key upper body landmarks (nose, shoulders, elbows, and wrists) and normalize their coordinates relative to the signing space, resulting in a compact 14-dimensional pose vector.
3.2 Self-Supervised Training of SHuBERT
Figure 4: Three masking strategies for sequence masking. Empirically, random masking results in the best translation performance (see Tab. 8).
Multi-Stream Input
We combine the four feature vectors described above—a 384-dimensional feature vector for each hand, a 384-dimensional face feature vector, and a 14-dimensional pose feature vector—resulting in 1,166 dimensions. We extract these features at every second frame.
Specifically, each channel first undergoes independent layer normalization to handle the varying scales of the different streams. Each channel’s normalized features are then projected through a separate linear layer to 256-dimension each. These projected channel-specific representations are concatenated to create a unified frame-level representation, which captures information from all modalities. This combined representation is then input to the transformer encoder to model temporal relationships across frames.
Sequence Masking
Inspired by prior work in self-supervised learning in speech [2, 20], SHuBERT learns by predicting masked elements of the input feature sequence, given the observed (unmasked) data. We use a masking strategy designed for multi-channel sign language input. We consider three types of masking, illustrated in Fig. 4: channel masking, which masks entire channels (e.g., all face and left hand features in a video) to learn cross-channel dependencies; time masking, which masks all channels at selected temporal positions (e.g., face, hand, and body pose in frames 20-40 in a given video); and random masking, which independently masks random frames in each channel. Based on our experiments comparing these strategies (Sec. 4.5), we ultimately chose random masking as our final approach. After replacing the masked inputs with a mask token, all of the feature frames are fed through the transformer encoder to produce contextualized representations.
Learning Objective
For the masked input elements, the representation at the last transformer layer is used to predict target pseudo-labels. Inspired by HuBERT [20], we use offline $k$-means clustering (separately for each of the four channels) to create discrete target pseudo-labels. Target labels are extracted independently for each channel, resulting in four separate cluster assignments per frame. A separate prediction head is used for each channel to predict the corresponding cluster assignment for masked regions.
The training objective is a cross-entropy loss between the target and predicted cluster assignments, computed over the masked positions.
This self-supervised training produces the pre-trained SHuBERT model, which can then be fine-tuned for downstream sign language tasks using appropriate task-specific prediction layers and losses.
4 Experiments and Results
Below we describe the experimental setup for self-supervised training of SHuBERT, followed by its adaptation as a foundation model for multiple sign language processing tasks: sign language translation (Sec. 4.2) isolated sign language recognition (Sec. 4.3), and phonological feature recognition (Sec. 4.4).
4.1 Pre-Training SHuBERT
Data and Pre-Processing
For pre-training SHuBERT, we use the Youtube-ASL dataset [55] excluding the clips that intersect with the OpenASL dataset [50], so that we can use OpenASL to evaluate SHuBERT. To maintain the same training set size as the original YouTube-ASL dataset, we replaced the removed content with ASL videos from YouTube-SL-25 [54] that are not present in YouTube-ASL. Our final pre-training dataset comprises approximately 987 hours of ASL content.
To reduce computation, we downsample the videos by removing every other frame The average frame rate of the post-processed videos is 14.89 fps.
Model Configuration and Training
We train a base model consisting of 12 transformer blocks, with each block having an embedding dimensionality of 768, feed-forward dimensionality of 3072, and 12 attention heads. The complete model contains 86M parameters. For each channel (face, left hand, right hand, and body posture), we use $k$-means on 10% of the data to create 256 discrete clusters that serve as prediction targets. Our masking strategy uses a span length of 3 frames (approximately 200ms), which is the average duration of fingerspelling a single letter in American Sign Language and is therefore roughly the smallest gesture length  [57]. We train the model for 400K steps using 8 NVIDIA A6000 (48GB) GPUs, with a total training time of approximately 7 days.
We optimize the model using Adam with a peak learning rate of $5\times 10^{-4}$, warming up for the first $8\%$ of updates followed by linear decay. Videos are batched to maintain efficiency while not exceeding 1,500 frames per GPU.
4.2 Sign Language Translation
For sign language translation, where the input sign language video is mapped to text in English, we use ByT5-Base [61] as a translation model to map from SHuBERT representations to written English, following prior work showing its strong performance on this task [52, 63]. We first extract video representations from SHuBERT and project them to ByT5’s input space through a simple linear layer. We train the combined model with a cross-entropy loss and label smoothing factor of 0.2. We use a two-phase training strategy. In the first phase, we train the translation system on the weakly labeled YouTube-ASL dataset (with OpenASL removed, as described above) for 250K steps. During this phase, we use the AdamW optimizer with a peak learning rate of $5\times 10^{-4}$ for ByT5 and a reduced learning rate of $5\times 10^{-5}$ for SHuBERT parameters (when fine-tuned). The learning rate follows a cosine schedule with 10K warmup steps. We use a batch size of 2 utterances per GPU with gradient accumulation over 8 steps and use weight decay of 0.1 for regularization.
In the second phase, we fine-tune on two target benchmark datasets (How2Sign [13] and OpenASL [50]) for 50K steps, using a lower learning rate of $10^{-4}$ and 5K warmup steps. We also evaluate in a zero-shot setting, without any additional fine-tuning, on a third dataset (for which no training data exists), FLEURS-ASL [53]. As often done when using pre-trained speech encoders [58], we use a learned weighted sum of features from all SHuBERT layers rather than using a single layer’s output. During decoding, we use beam search with a beam width of 5, and a maximum sequence length of 384 tokens.
Table 1: Translation results on How2Sign. SSL: self-supervised learning (yes/no); PT: pre-training data.
[…]
Private data
Tanzer [52] $\times$ $\sim$2,800 18.1 50.8
Zhang et al. [63] $\times$ $\sim$6,600 21.1 55.7
Publicly available data
SSVP [44] $\checkmark$ 1,054 15.5 49.6
Uthus et al. [55] $\times$ 984 12.4 46.6
Tanzer and Zhang [54] $\times$ 3,207 15.4 47.9
SignMusketeers [19] $\checkmark$ 984 14.3 -
Ours $\checkmark$ 984 16.2 49.9
Table 2: Translation results on OpenASL
[…]
Publicly available data
OpenASL[50] $\times$ - 6.7 31.1
GloFE-VN [33] $\times$ - 7.1 36.7
C2RL [10] $\times$ - 13.2 -
Ours $\checkmark$ 984 23.2 60.6
Table 3: Translation results on FLEURS-ASL
[…]
Private data
Tanzer [52] $\times$ $\sim$2,800 5.8 45.4
Publicly available data
Tanzer [53] $\times$ 984 3.9 38.3
Tanzer and Zhang [54] $\times$ 3,207 4.4 40.1
Ours $\checkmark$ 984 4.7 41.2
We evaluate the final model on the three benchmark test sets. Tab. 1, Tab. 2, and Tab. 3 show translation results on How2Sign, OpenASL, and FLEURS-ASL, respectively. We evaluate using the standard BLEU[39, 40] and BLEURT[47] translation metrics.
In all cases, our results using SHuBERT improve over the best prior published results using publicly available training data. On How2Sign, SHuBERT achieves a +0.7 BLEU/+0.3 BLEURT improvement over the prior state-of-the-art result (using public data), belonging to SSVP which used slighty more pre-training data (1,054 vs. 984 hours, as they included the training data of How2Sign in pre-training). Better published results exist (as shown in Tab. 1), but they use private fine-tuning datasets so we cannot reproduce their settings nor compare to them meaningfully.
In the case of OpenASL, SHuBERT’s improvement over the best prior result is particularly large (+10.0 BLEU/+23.9 BLEURT). This may be attributable to pre-training on similar-domain (native, natural rather than interpreted signing) data in YouTube-ASL (but, as previously mentioned, not any of the same data). The distinction between interpreted signing in a constrained visual environment (as in How2Sign and FLEURS-ASL) and natural signing in a less constrained environment (as in OpenASL and much of YouTube-ASL) is an important one that has not received sufficient attention in the sign language understanding literature, and is worth exploring further in future work. The properties of the visual environment affect the difficulty of the task, and natural signing has different characteristics from those of interpreted sign language, often described as “translationese” [16, 29].
On FLEURS-ASL, a dataset designed for testing only, SHuBERT demonstrates strong performance in a zero-shot setting, surpassing both prior methods, which used over 3x as much pre-training data (3,207 hours) as ours.
We also note that most previous approaches use additional enhancement techniques shown to contribute to their final scores, such as an additional contrastive learning with labelled data [44], joint training with text machine translation [63], and multi-tasking with random dynamic clips [53]. In contrast, our setup uses none of those. It is possible that incorporating such techniques into our framework will further improve scores, but we leave this for future work.
4.3 Isolated Sign Language Recognition
For isolated sign language recognition (ISLR), the task of classifying a short video of a single sign, we compare a range of approaches for using SHuBERT, including learning LoRA adapters [21], fully fine-tuning, and joint multi-task training on multiple ISLR datasets. Unless otherwise specified, for all experiments, we train for 125 epochs with a batch size of 128 and perform early stopping according to validation results. We use an Adam optimizer [28] with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. For adapting SHuBERT to classification tasks, we first average SHuBERT representations across the time dimension and add a batch-norm layer followed by a linear layer as the classification head.
For LoRA training, we learn a rank-1 LoRA module for each linear layer in SHuBERT while keeping all the other weights frozen, increasing by only 0.197% the number of parameters. In addition to the aforementioned hyperparameters, we reduce the learning rate of the LoRA modules to 1/10 of the classification head’s and also use a 0.1 label smoothing for further regularization.
For the case of full fine-tuning of SHuBERT, the training protocol is largely the same as for LoRA training, except that we re-initialize the last layer of SHuBERT which we find to give a small performance improvement.
Finally, inspired by prior work such as [55] and [44], where training on larger datasets greatly enhances the performance on smaller ones, we fine-tune SHuBERT jointly on three ISLR datasets: ASL Citizen [11], WLASL2000 [32, 37] (cleaned version) and Sem-Lex [26]. Since these datasets have different labels, we frame the learning as multi-task training; each task has its own classifier, with the fine-tuned SHuBERT encoder shared across tasks. The losses from the three datasets are weighted equally. The hyperparameters are the same as in full fine-tuning except for the batch size, which is reduced to 100 due to GPU memory constraints.
Our ISLR results on ASL Citizen, WLASL2000 (clean) and Sem-Lex are shown in Tab. 4, Tab. 5 and Tab. 6, respectively. We report Recall at 1, 5 and 10. We achieve state-of-the-art performance on all datasets except WLASL2000 (clean). [^2] Note: As described in the caption of Tab. 6, the Sem-Lex results in [26] are not comparable with other methods, including ours. As shown in all three tables, joint training outperforms the other two training strategies on all datasets we experiment on, with a margin of 3.8% on average. Rank-1 LoRA, although it is the worst performer of our results on two of the three datasets, still outperforms previous SOTA results on ASL Citizen and Sem-Lex. For WLASL2000 (clean), Ours (joint) yields competitive results that are nearly on par with MaskFeat, a model trained specifically for ISLR.
Table 4: ISLR accuracy results on the ASL Citizen dataset.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
ST-GCN[11] 0.60 0.82 0.88
SignCLIP[25] 0.60 0.84 0.89
I3D[11] 0.63 0.86 0.91
Ours (rank=1 LoRA) 0.65 0.87 0.91
Ours (fine-tune) 0.65 0.87 0.91
Ours (joint) 0.68 0.90 0.93
Table 5: ISLR accuracy results on the WLASL2000 (clean) dataset.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
SVT[45] 0.19 - -
BEVT[45] 0.61 - -
VideoMAE[45] 0.69 - -
MaskFeat[45] 0.79 - -
Ours (rank=1 LoRA) 0.68 0.89 0.94
Ours (fine-tune) 0.71 0.92 0.95
Ours (joint) 0.76 0.94 0.96
Table 6: ISLR accuracy results on the Sem-Lex dataset. The result marked with an asterisk () is not directly comparable to Ours as it is for a reduced (and easier) test set, as mentioned in [25]. Additionally, the dataset version released by [26] has a significant fraction of videos missing.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
Sem-Lex[26] $0.69^{*}$
SignCLIP[25] 0.30 0.48 0.55
Ours (rank=1 LoRA) 0.54 0.74 0.80
Ours (fine-tune) 0.53 0.73 0.79
Ours (joint) 0.58 0.78 0.83
4.4 Phonological Feature Recognition
Table 7: Phonological feature recognition accuracy results of two datasets, Sem-Lex and ASL Citizen.
Phonological Feature Rec@1$\uparrow$
Sem-Lex ASL Citizen
Major Location 0.8477 0.9022
Minor Location 0.7130 0.8000
Second Minor Location 0.7328 0.8118
Contact 0.8684 0.9157
Thumb Contact 0.8474 0.8752
Sign Type 0.8464 0.9154
Repeated Movement 0.8265 0.8993
Path Movement 0.7275 0.7942
Wrist Twist 0.9058 0.9300
Selected Fingers 0.7953 0.8344
Thumb Position 0.8604 0.8819
Flexion 0.7264 0.7773
Spread 0.7942 0.8480
Spread Change 0.8160 0.8658
Nondominant Handshape 0.7632 0.8432
Handshape 0.6293 0.7080
Average 0.7938 0.8502
We also conduct experiments on phonological feature recognition, that is the classification of linguistic features of signs, for two of the ISLR datasets. We report the Recall at 1 (prediction accuracy) for 16 commonly used phonological features (from ASL-LEX 2.0 [46]) in Tab. 7. The training setups and hyperparameters are identical to those of the full fine-tuning method in Sec. 4.3, except that we now train 16 classification heads simultaneously. We also find that removing weight decay gives a slight performance boost.
To the best of our knowledge, no previous work has reported phonological feature recognition accuracies on the ASL Citizen dataset. Similarly, though the Sem-Lex authors [26] report phonological feature prediction accuracies, they are not comparable to ours, which are computed on the entire test set available to the public. Thus, we hope that our results in Tab. 7 can serve as a benchmark for future work.
4.5 Ablations
We conduct ablations, using the How2Sign translation task, to measure the importance of different factors in the SHuBERT pre-training and adaptation to the downstream task. Unless stated otherwise, in the ablation experiments we pre-train SHuBERT for 100K steps (instead of the full 400K) for a faster turnaround and freeze SHuBERT for the downstream task.
Masking Strategies
Table 8: Comparison of masking strategies in SHuBERT pre-training. SHuBERT is frozen and stopped after 100K steps, and fine-tuned and evaluated on How2Sign only.
Masking strategy BLEU-1 BLEU BLEURT
Channel masking 14.5 2.6 29.9
Time masking 15.1 2.3 31.2
Random masking 15.4 2.2 31.4
Tab. 8 compares translation performance when using each of the three masking strategies (Fig. 4). We observe different signals from different metrics. While the BLEU scores suggest channel masking to be the best, random masking produces better BLEURT scores. Contradicting signals between different evaluation metrics for sign language translation has also been observed in prior work [63]. We chose to priotize BLEURT, as it generally has better alignment with human judgements of translation quality [15], and therefore use random masking in all of our other experiments.
Data Scaling Behavior of SHuBERT
Table 9: Impact of pre-training data size on SHuBERT’s performance.
Hours of pre-train data BLEU-1 BLEU BLEURT
984 15.4 2.2 31.4
98 12.7 0.7 29.1
In addition to the full pre-training dataset, we also train SHuBERT from scratch on a randomly selected 10% of the full data. In Tab. 9 we see that there is a noticeable drop in performance in BLEU and BLEURT, suggesting that data size is important. With multiple larger datasets now available—BOBSL[1], JWSign[18] and YouTube-SL-25[54] contain approximately 1500, 2500, and 3200 hours of data respectively—we expect that expanding the SHuBERT pre-training data may further improve performance. In addition, these larger datasets also include more language diversity, which may also improve performance and/or applicability to additional languages.
Isolating SHuBERT’s Impact on Performance
Table 10: Contribution of several components of SHuBERT: direct use of input video features (None) vs. using SHuBERT’s last layer vs. weighted combination of all layers. In all cases SHuBERT is frozen.
Layer of SHuBERT BLEU-1 BLEU BLEURT
None 15.3 2.5 31.6
Last Layer 21.4 4.7 35.0
Weighted Sum 29.3 7.1 39.5
To understand SHuBERT’s impact on translation performance, we conduct three experiments, shown in Tab. 10. In our baseline experiment (“None”), we directly feed the projected 4-channel features (face, left hand, right hand, body pose) to the ByT5 translation model, bypassing SHuBERT entirely, resulting in fairly poor performance. When we instead pass these features through a frozen pre-trained SHuBERT (400k steps, random masking) and use its final layer’s output (“Last Layer”), we see significant improvement. Finally, computing a learned weighted combination of all SHuBERT layers (“Weighted Sum”) further improves performance. These results demonstrate that each component of SHuBERT contribute to translation performance.
Frozen vs. Fine-Tuned SHuBERT
Table 11: Effect of fine-tuning on translation performance: Fine-tuning SHuBERT along with the translation model ($\checkmark$) vs. using frozen SHuBERT representations ($\times$).
Fine-tune SHuBERT BLEU-1 BLEU BLEURT
$\times$ 21.4 4.7 35.0
$\checkmark$ 30.0 7.5 39.9
We also investigate the impact of fine-tuning during translation training, as shown in Tab. 11. We compared two scenarios: using a frozen SHuBERT (400k steps, random masking) and only fine-tuning ByT5, versus fine-tuning both SHuBERT (from the same base model) and ByT5 together. Both scenarios use features from SHuBERT’s last layer. Fine-tuning SHuBERT leads to substantial improvements compared to keeping it frozen, when using the final layer. However, referring back to Tab. 10, the relatively small difference between the fine-tuned performance and that of the frozen and layer-weighted SHuBERT is noteworthy. This observation is promising for low-resource sign languages, where we may have plentiful unlabeled video data for pre-training, but very limited parallel data for translation training.
5 Conclusion
SHuBERT, our proposed self-supervised approach for learning sign language video representations, yields a transformer encoder that maps from multiple feature streams (face and hand appearance and upper body pose) to a stream of per-frame contextual representations. A single base SHuBERT model, when adapted to a range of sign language processing benchmarks including both translation and isolated sign recognition, achieves strong performance on all of them and almost always improves over the state of the art. SHuBERT is trained on public data and will be released upon publication. Based on its strong performance on the tasks studied here, we expect that SHuBERT can serve as a base model for a broad range of sign language processing tasks. In addition to the directions for future work mentioned in the relevant experiment sections above, SHuBERT has the potential for further improvement using techniques like iterative re-clustering and re-training, such as those used for similar speech models [20].
Limitations
Our work has several limitations. First, although the results are competitive with or outperform prior work, the absolute performance is still quite poor. Neither our model nor others can replace human interpreters for broad-domain sign language translation. Second, our training data volume is significantly smaller than that of typical self-supervised speech and text models. We cannot say with certainty how our model would scale up to much larger datasets. Finally, the current scope of our work is limited to American Sign Language. Although sign languages use the same channels and share many elements, we do not know how well our model would generalize to other languages. Future work could address these limitations by expanding the training dataset and training on data from other sign languages.
Acknowledgment
We are grateful to Shiry Ginosar, Anand Bhattad, Ju-Chieh Chou, and Chung-Ming Chien for their valuable suggestions throughout this project.
Supplementary Material
6 ASL Phonological Feature Classification Details
American Sign Language (ASL) can be described through a set of phonological features, similarly to the description of spoken languages via features. These features capture the essential components of sign formation, including hand configuration, movement patterns, and spatial relationships. Tab. 12 presents the number of classes for each of the phonological features we use in our phonological feature prediction analysis for Sem-lex and ASL-Citizen, and below we list the values of each feature. This commonly used feature set is from ASL-LEX 2.0 [46]).
Table 12: Number of classes for each phonological feature represented in two ASL datasets, Sem-Lex and ASL Citizen. Most features have the same number of classes across datasets, while a few features have values that don’t appear in one of the datasets (for example, Second Minor Location has 37 classes that appear in Sem-Lex and 38 classes in ASL Citizen).
Phonological Feature Number of Classes
Sem-Lex ASL Citizen
Major Location 5 6
Minor Location 37 37
Second Minor Location 37 38
Contact 2 2
Thumb Contact 3 3
Sign Type 6 6
Repeated Movement 2 2
Path Movement 8 8
Wrist Twist 2 2
Selected Fingers 12 12
Thumb Position 2 2
Flexion 8 8
Spread 3 3
Spread Change 3 3
Nondominant Handshape 56 57
Handshape 58 58
Handshape
v, 5, y, h, open_b, c, baby_o, flat_h, o, l, 1, a, open_8, w, curved_5, d, flatspread_5, i, f, s, p, flat_b, curved_4, flat_o, g, open_e, 4, closed_b, bent_1, 3, flat_horns, goody_goody, flat_m, bent_v, flat_1, r, 8, curved_v, open_h, curved_1, horns, flat_ily, flat_n, bent_l, stacked_5, ily, e, flat_v, curved_l, spread_open_e, curved_h, 7, closed_e, t, flat_4, open_f, k, and spread_e.
Nondominant Handshape
v, 5, y, none, open_b, Dominance Condition Violation, B, 1, a, open_8, C, s, h, o, flat_b, curved_5, p, c, S, closed_b, 4, flat_m, bent_v, flat_1, flat_h, baby_o, curved_v, i, f, bent_1, Symmetry Violation, flatspread_5, flat_o, curved_1, open_h, stacked_5, g, l, bent_l, 3, 8, spread_open_e, e, horns, w, r, Lax, curved_l, open_e, flat_4, O, curved_b, A, ily, flat_v, and flat_horns.
Minor Location
Neutral, Head Away, Body Away, Hand Away, Palm, Finger Tip, Forehead, Finger Front, Mouth, Chin, Other, Upper Arm, Torso Top, Forearm Back, Cheek Nose, Wrist Front, Palm Back, Finger Back, Finger Radial, Under Chin, Finger Ulnar, Wrist Back, Shoulder, Arm Away, Forearm Ulnar, Torso Mid, Heel, Clavicle, Eye, Forearm Front, Neck, Torso Bottom, Upper Lip, Head Top, Elbow Back, Hips, and Waist.
Second Minor Location
Neutral, Head Away, Torso Bottom, Finger Tip, Hand Away, none, Palm, Forearm Back, Finger Back, Body Away, Torso Top, Finger Front, Chin, Arm Away, Upper Arm, Finger Ulnar, Eye, Hips, Neck, Palm Back, Forearm Front, Finger Radial, Mouth, Heel, Torso Mid, Other, Waist, Cheek Nose, Forehead, Elbow Back, Under Chin, Clavicle, Shoulder, Forearm Ulnar, Head Top, Upper Lip, and Forearm Radial.
Sign Type
Symmetrical Or Alternating, One Handed, Dominance Violation, Asymmetrical Different Handshape, Asymmetrical Same Handshape, and Symmetry Violation.
Path Movement
Curved, Back And Forth, Straight, Circular, None, Z-shaped, Other, and X-shaped.
Flexion
Fully Open, Curved, Bent, Flat, none, Fully Closed, Stacked, and Crossed.
Selected Fingers
im, imrp, p, i, t, m, ip, imp, mr, imr, r, and mrp.
Major Location
Neutral, Head, Body, Hand, and Arm.
Flexion Change
1.0, 0.0, and none.
Spread Change
1.0, 0.0, and none.
Thumb Contact
1.0, 0.0, and none.
Spread
1.0, 0.0, and none.
Thumb Position
Closed and Open.
Repeated Movement
1.0 and 0.0.
Contact
1.0 and 0.0.
Wrist Twist
0.0 and 1.0.
7 Cluster Samples
We visualize clustering results for the face, left hand, right hand, and upper body pose in Figs. 5, 6, 7 and 8. All cluster samples were randomly selected (i.e., without manual curation or cherry-picking). Each row represents a cluster. For each channel (Face, Left hand, Right hand, Upper Body), we include 10 random examples for 5 random clusters. While there is variability within each cluster, and some clusters contain a large mix of poses, we can also see a great deal of systematic behavior, where the images in a cluster tend to correspond to similar gestures regardless of signer appearance or other visual properties. The caption for each figure provides our interpretations of some of the clusters.
Figure 5: Sample face clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 141 includes mainly open-mouthed expressions with raised eyebrows, cluster 57 seems to capture closed or squinting eyes with neutral mouths, and cluster 251 corresponds to a slightly tilted head with direct gaze and little mouth opening. NOTE: For clarity, we show unblurred cropped faces here.
Figure 6: Sample right hand clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 42 captures mainly multi-finger pointing gestures. Cluster 76 corresponds to clasped or overlapped hands in resting positions.
Figure 7: Sample upper body clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 5 seems to correspond to configurations of the upper body pose where the right hand is at shoulder level and the left hand is down. Cluster 43 seems to be a configuration where the two hands are raised and close to each other near the chest, and the signer is facing slightly to the right. This might correspond to signs being performed with both hands being involved/active. Cluster 36 appears similar, but with the hands slightly farther apart. Cluster 19 is generally similar to cluster 43, except that the signer tends to be facing slightly to the left. Finally in cluster 168, the right hand is usually above the shoulder and close to the face/head.
Figure 8: Sample left hand clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 125 shows pointing configurations with the index finger extended. Cluster 115 generally corresponds to closed fist formations oriented with the thumb on top. Cluster 51 seems to be a mix of poses without a consistent description. Cluster 98 seems to include mainly transitional hand movements around the chest.
8 Translation examples
We provide example translations produced by our model fiven inputs from three ASL datasets: How2Sign (instructional content) Tab. 13, FLEURS-ASL (zero-shot setting) Tab. 14, and OpenASL (general domain with native signers and varying background) Tab. 15.
Table 13: Qualitative translation examples from the How2Sign dataset, comparing SHuBERT-based translations to previous models.
(1) Reference And that’s a great vital point technique for women’s self defense.
Uthus et al. [55] It’s really great for women’s self defense.
Rust et al. [44] This is a really great point for women’s self defense.
Ours If you’re a bigger person we’re talking about really self defense here.
(2) Reference In this clip I’m going to show you how to tape your cables down.
Uthus et al. [55] In this clip we’re going to show you how to cut a piece of clay.
Rust et al. [44] In this clip I’m going to show you how to clip the cable, the cable.
Ours In this clip I’m going to show you how to brand out the cable strings.
(3) Reference In this segment we’re going to talk about how to load your still for distillation of lavender essential oil.
Uthus et al. [55] In this clip we’re going to talk about how to feed a set of baiting lizards for a lava field oil.
Rust et al. [44] In this clip we’re going to talk about how to feed the trail for draining clean for laborer oil.
Ours In this clip we’re going to talk about how to take our stick for disinfectant oil.
(4) Reference You are dancing, and now you are going to need the veil and you are going to just grab the veil as far as possible.
Uthus et al. [55] Their hopping and dancing is now, they’re going to need their squat and squat and they’re going to be able to move independently.
Rust et al. [44] So that she’s going to get her hips up as far as she can, and now she’s going to lift her head up as far as possible.
Ours Her dancing and now she needs her feather to grab it with her foot as far as possible.
(5) Reference But if you have to setup a new campfire, there’s two ways to do it in a very low impact; one is with a mound fire, which we should in the campfire segment earlier and the other way to setup a low impact campfire is to have a fire pan, which is just a steel pan like the top of a trash can.
Uthus et al. [55] But if you have to set up a new campfire, there are two ways to do a low impact fire, one is a cone fire, which we have to do in the tent earlier, and the other one is to set up a campfire in a fire pan.
Rust et al. [44] But if you have to set up a new campfire, this is one way to do it in a low impact. One is a monk fire. One is a campfire. The other one is to set a campfire in a campfire. That’s just a post like the top of the post.
Ours But if you have to set a new campfire, there are two ways to do a low impact one is a bond fire, which we should do in your campfire, another one is to set a campfire in a fire pan that is just just set a pan like the top of it pan.
(6) Reference So, this is a very important part of the process.
Uthus et al. [55] Alright, let’s get started.
Rust et al. [44] It’s an important part of the process.
Ours This is a very important part of the process.
Table 14: Qualitative translation examples from the FLEURS-ASL dataset, comparing SHuBERT-based translation (zero-shot) to a previous approach [53].
(1) Reference During the 1980s he worked on shows such as Taxi, Cheers, and The Tracy Ullman Show.
Tanzer [53] In the 1980s, she worked in theaters like taxesi, cheesy, and tracy.
Ours In the 1980’s, she worked for theaters like Taxi Chers, Tracy Ullman Shaw.
(2) Reference The rise of new technologies allows us to see and investigate brain structures and processes never seen before.
Tanzer [53] There is a new technique to detect brains and vision.
Ours Increasing new technology that allows people to consider investigating their brain structures and brain structures.
(3) Reference The Articles required unanimous consent from all the states before they could be amended and states took the central government so lightly that their representatives were often absent.
Tanzer [53] The law requires all states to agree on a standard and that it is a legal requirement.
Ours The state’s agreement requires all standardized agreements to remove the standards of representatives from the state to represent the state’s amendments.
Table 15: Qualitative translation examples from the OpenASL dataset, comparing SHuBERT-based translations to a previous model [50]
. (1) Reference thank you Shi et al. [50] thank you Ours thank you (2) Reference come on Shi et al. [50] come on Ours maybe (3) Reference now i’ve come this far and it ’s a different team Shi et al. [50] how do you feel about it Ours it feels like a crash in the team (4) Reference i was there from the beginning to the end and time went by fast Shi et al. [50] the students were thrilled by this Ours i just wanted to leave because it went ahead and started (5) Reference i’m here at nad’s 50th wow Shi et al. [50] the nad has been ¡unk¿ for many years Ours well the nad is shocked to have 50 years of dhh (6) Reference i entered the yap 2018 competition and won Shi et al. [50] the competition was started with ideas Ours i enrolled in that competition in 2018 and then i won (7) Reference you can check out their kickstarter in the link below Shi et al. [50] you can watch the conversation at lake county Ours you can check out their kickstarter link below (8) Reference that is one thing i found interesting and wanted to share with you today Shi et al. [50] i also am the president of the jr. nad conference here Ours that’s one interesting thing she wanted to share with you (9) Reference those are the different types of bills Shi et al. [50] schools have switched to teaching students Ours i looked at several different types of interpreting services (10) Reference dry january has picked up in popularity since it began in 2012 Shi et al. [50] krispy kreme is bringing back its original playstation in 2016 Ours the qury dry january started in 2012 (11) Reference we will be happy to respond give you support and listen to your concerns Shi et al. [50] please review and submit your time passion and support this important issue Ours the nad is willing to respond and support your concerns (12) Reference there were videos posted on the internet that showed a person walking on the grass completely engulfed in flames Shi et al. [50] a video shows the officer walking up to his shoulder and before he was shot Ours videos posted on social media of him walking on a grass walking completely with fire (13) Reference and people would become carpenters laborers mechanics plowers and farmers Shi et al. [50] the next year 1880 the nad was established in the first operation 30 of the house in 2015 Ours and he was forced to wear a wearing a labover meganic and a financial warper (14) Reference for nad youth programs related information please contact us via facebook at the nad youth programs or email us through Shi et al. [50] you can contact us through our website where you can check our facebook page online at ¡unk¿ Ours if you want to contact the nad youth program you can contact us through our facebook page at the nad youth program or youth program through our website (15) Reference last week suspects gregory mcmichael and his son travis were arrested and charged with felony murder and aggravated assault Shi et al. [50] last week a black man named ¡unk¿ ¡unk¿ was arrested and charged with felony murder and aggravated assault Ours last week two suspects gregory mcmichael and his son travis were arrested and charged with felony murder and wounded by another gravated assaulter